OpenStudy (anonymous):

In Iman's large aquarium the ratio of goldfish to angelfish is 3:2. If there are 115 fish in his aquarium hiw many more goldfish are there then angelfish?

OpenStudy (anonymous):

ratio of goldfish to angelfish is 3:2 = let 3x be no of goldfish 2x be no of angelfish givn here are 115 fish in his aquarium 3x+2x=115 can u proceed find x val

OpenStudy (anonymous):

Are sure this is how you do it then thanks big thanks

OpenStudy (anonymous):

yes its procedure

OpenStudy (anonymous):

So there are 23 Moore

OpenStudy (anonymous):

So 5x= 115 X=23 Correct?

OpenStudy (anonymous):

yes nw find no of goldfish=3x no of angelfish=2x

OpenStudy (anonymous):

Howdo you do that

OpenStudy (anonymous):

Plz help

OpenStudy (anonymous):

Oh I see you substitute

OpenStudy (anonymous):

no of goldfish=3x=3*23 =?.... no of angelfish=2x=2*23 =?...... find them

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

giv me medal if u lik

OpenStudy (anonymous):

I d, how to do that but if I did I would give you one